OpenStudy (anonymous):

Find the derivative of the function g(x)=(sqrt3-x) directly from the definition. (For this question, are we using the limits?)

OpenStudy (anonymous):

is it sqrt(3-x)?

OpenStudy (anonymous):

\[g(x)=\sqrt{3-x}\]

OpenStudy (jdoe0001):

yeap as h -> 0

OpenStudy (anonymous):

okay..lets get to work

OpenStudy (anonymous):

So far I got this: \[\lim_{h \rightarrow 0}\frac{ \sqrt{3-(x+h)} -\sqrt{3-x} }{ h }\]

OpenStudy (jdoe0001):

I've gotten that much as well

OpenStudy (jdoe0001):

hmmm I don't see it coming out cleanly

OpenStudy (anonymous):

hmmmm...

OpenStudy (solomonzelman):

I am hanging on other question still so I will see how it goes.

OpenStudy (anonymous):

hmmm.....

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\lim_{h \rightarrow ~0}\frac{\sqrt{3-(x+h)}-\sqrt{3-x}}{h}}\) my first thought is to rationalize the numerator

OpenStudy (solomonzelman):

for any (a-b) conjugate is (a+b).....

OpenStudy (solomonzelman):

you don't want me to do this problem for you, do you ?

OpenStudy (anonymous):

I did try rationalizing the numerator, but I'm not getting it correctly. If you don't mind, can you show me how?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \large\color{slate}{\displaystyle\lim_{h \rightarrow ~0}\frac{\sqrt{3-(x+h)}-\sqrt{3-x}}{h}} }\) \(\large\color{black}{ \displaystyle \large\color{slate}{\displaystyle\lim_{h \rightarrow ~0}\frac{ \color{red}{(}\sqrt{3-(x+h)}-\sqrt{3-x}\color{red}{)~\times (\sqrt{3-(x+h)}+\sqrt{3-x})} }{h\color{red}{~\times (\sqrt{3-(x+h)}+\sqrt{3-x})} }} }\)

OpenStudy (solomonzelman):

see how I am multiplying times a conjugate on top and bottom ?

OpenStudy (anonymous):

yes..

OpenStudy (solomonzelman):

can you simplify the top for me please?

OpenStudy (solomonzelman):

what if it was [ sqrt(A)-sqrt(B) ] * [ sqrt(A) + sqrt(B) ] , what would you then tell me ?

OpenStudy (anonymous):

The square roots get cancel out

OpenStudy (solomonzelman):

and what would you then be getting ?

OpenStudy (anonymous):

(A-B)(A+B)

OpenStudy (solomonzelman):

please try again

OpenStudy (solomonzelman):

(g-h)(g+h)=g^2-h^2 is this correct ?

OpenStudy (anonymous):

No....

OpenStudy (solomonzelman):

that is a rule, called a difference of squares.

OpenStudy (solomonzelman):

you can prove why it is true, simply be expanding and canceling out the terms... you will end up with g^2-h^2.

OpenStudy (anonymous):

Oh yeah... Sorry I forgot

OpenStudy (solomonzelman):

so then, \(\large\color{black}{ \displaystyle (\sqrt{A}+\sqrt{B})(\sqrt{A}-\sqrt{B})=(\sqrt{A~})^2-(\sqrt{B~})^2}\)

OpenStudy (solomonzelman):

am I correct ?

OpenStudy (anonymous):

Yes. final anwer is A-B

OpenStudy (solomonzelman):

yes, so \(\large\color{black}{ \displaystyle (\sqrt{a}-\sqrt{b})(\sqrt{a}+\sqrt{b})=a-b }\)

OpenStudy (solomonzelman):

Now, look at this (where all this discussion about difference of squares started). \(\large\color{black}{ \displaystyle \large\color{slate}{\displaystyle\lim_{h \rightarrow ~0}\frac{ \color{red}{(}\sqrt{3-(x+h)}-\sqrt{3-x}\color{red}{)~\times (\sqrt{3-(x+h)}+\sqrt{3-x})} }{h\color{red}{~\times (\sqrt{3-(x+h)}+\sqrt{3-x})} }} }\)

OpenStudy (solomonzelman):

do you see this \(\large\color{black}{ \displaystyle (\sqrt{a}-\sqrt{b})(\sqrt{a}+\sqrt{b}) }\) form in the numerator ?

OpenStudy (anonymous):

Yes. It would be 3-(x+h)-3-x?

OpenStudy (solomonzelman):

yes, there you go.... and that simplifies to what ?

OpenStudy (solomonzelman):

oh, wait you are wronf

OpenStudy (solomonzelman):

3-(x+h)-(3-x) like this.

OpenStudy (solomonzelman):

see why is it like this with the parenthesis ?

OpenStudy (solomonzelman):

you are subtracting the entire b, and your entire b is 3-x, (not just x or just 3) so you are subtracting the (3-x). and that is why we get a - b (3-(x+h)) - (3-x)

OpenStudy (solomonzelman):

understand why we need to write in parenthesis 3-(x+h)-(3-x) ?

OpenStudy (anonymous):

Yes... We need to distribute negative sign

OpenStudy (anonymous):

@SolomonZelman .....wowww...I GOT IT..:)...cool man.. it gets simplified to -h... h and h get cancelled... apply de limits.... the answer comes...:)

OpenStudy (solomonzelman):

yes, sidarth, then you plug in zero for h (which is valid from that point - i.e. without making denom. undefined)

OpenStudy (solomonzelman):

3-(x+h)-(3-x) simplifies to ?

OpenStudy (anonymous):

@SolomonZelman ...yeah...:)

OpenStudy (anonymous):

Oh! I got the final answer! \[\frac{ -1 }{ 2\sqrt{3-x} }\]

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \large\color{slate}{\displaystyle\lim_{h \rightarrow ~0}\frac{ \color{red}{(}\sqrt{3-(x+h)}-\sqrt{3-x}\color{red}{)~\times (\sqrt{3-(x+h)}+\sqrt{3-x})} }{h\color{red}{~\times (\sqrt{3-(x+h)}+\sqrt{3-x})} }} }\) \(\large\color{black}{ \displaystyle \large\color{slate}{\displaystyle\lim_{h \rightarrow ~0}\frac{ -h }{h\color{red}{~\times (\sqrt{3-(x+h)}+\sqrt{3-x})} }} }\) \(\large\color{black}{ \displaystyle \large\color{slate}{\displaystyle\lim_{h \rightarrow ~0}\frac{ -1 }{\sqrt{3-(x+h)}+\sqrt{3-x} }} }\) \(\large\color{black}{ \displaystyle =\large\color{slate}{\displaystyle\frac{ -1 }{\sqrt{3-(x+0)}+\sqrt{3-x} }} }\) \(\large\color{black}{ \displaystyle =\large\color{slate}{\displaystyle\frac{ -1 }{\sqrt{3-x}+\sqrt{3-x} }} }\) \(\large\color{black}{ \displaystyle =\large\color{slate}{\displaystyle\frac{ -1 }{2\sqrt{3-x} }} }\)

OpenStudy (solomonzelman):

coirrect

OpenStudy (solomonzelman):

(( Same way you would normally differentiate and do the chain rule for "3-x". ))

OpenStudy (anonymous):

oh okay! Thanks for explaining it to me! ^_^

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{d}{dx}\left[ \sqrt{3-x}~\right] ~=~(3-x)^{1/2 \color{red}{~~-1}}~\color{blue}{ \times \frac{d}{dx}\left[ 3-x\right]} }\) \(\large\color{black}{ \displaystyle \frac{d}{dx}\left[ \sqrt{3-x}~\right] ~=~(3-x)^{-1/2 }~\color{blue}{ \times (-1)} }\) \(\large\color{green}{ \displaystyle \frac{d}{dx}\left[ \sqrt{3-x}~\right] ~=~\frac{-1}{\sqrt{3-x}}}\)

OpenStudy (solomonzelman):

in red I decremented the power by 1, and blue is the chain rule for the part inside the square root.

OpenStudy (solomonzelman):

yw

OpenStudy (solomonzelman):

do you know derivatives though, have you learned them ?

OpenStudy (solomonzelman):

anyway... have fun (I guess)